If tan x = -and x is in the fourth quadrant, find the following values. (a) sin (1) (b) cos (3) (c) tan(2x)
Answer #1

Given tanx = = opposite adjacent & (y)? (4) + (3) I XO ya = 16+9 y = $25 25 sinx= opposite & 4/5 hypotenuse I Given x is in 4A negative x/2 lies in pond quadrant, coline is cos (1/2) = -2 l tan 2x = 2 tanx = 2(-4) 7-19) 1- tanx Therefore, sin (2/2) =